# Shifting the Lens: Detecting Malicious npm Packages using Large Language Models

### v2

@@ III-D3 Step 3: Final Report
The final report prompt (Table III) is designed to select and refine the best report from the critical analysis step. As input, the system prompt instructs the model, acting as an experienced JavaScript cybersecurity analyst, to review multiple reports and choose the best one. The task involves eliminating irrelevant signals by reviewing the prior generated reports (section III-D2) and setting the malware score to 1 for malicious code to emphasize consistency between the final score and the report content. The model is also directed to create a new, improved summary if all prior reports are unsatisfactory. The step also involves sequentially feeding the model system role prompts from Step 1 (Section III-D1) to remind the objectives and ensure a thorough code reevaluation. For the user role prompt, we feed the previously generated response report from section III-D2 and the file to be evaluated. The model uses system role prompts, user inputs, and preceding critical response reports to output a conclusive summary report, operating at a lower temperature of 0.5 and a $top_{p}$ value of 0.5 to enhance focus and precision. We include the prior model-generated responses and the file as a user prompt for the Critical Reports and Final Report step.
III-D4 Step 4: Report Evaluation
@@
Table IV contains the shell scripts of real-world malicious packages “corporate-delegate-packages” and the associated GPT-4-generated final report. We collect the JSON-formatted final report for each file, which includes the purpose of the code, source, sink, flows, anomalies, analysis, rationality behind model-derived conclusions, and scores for confidence, obfuscation, malware, and security risk. Reviewers can prioritize evaluations based on the malware score. Based on our scoring rules, a package is prioritized for human evaluation if any file has a malware score exceeding 0.5 (Section IV-A2). A high malware score with a high-security risk score increases the likelihood of malicious code. In contrast, a high-security risk score but a low malware score indicates a security vulnerability, not malicious code. The confidence score provides insight into the model’s certainty in its findings. An obfuscated score is helpful to understand whether the attacker used obfuscation to hide the code intent. Reviewers can review the purpose and conclusion parts (R1 & R7 in Table  IV) to comprehend the code’s intent. If further verification is needed, they can analyze the code review metrics (R2-R6 in Table IV).
@@ IV-A1 CodeQL Static Analysis
Prior studies [14, 34, 17, 30, 42, 16] have evaluated malicious packages to identify relevant features to detect malicious packages. Froh et al., in a recent 2023 study [16], introduced a differential static analyzer leveraging custom CodeQL queries to identify JavaScript malicious code. CodeQL is a static analyzer with a semantic code analysis engine from GitHub that detects vulnerabilities across different programming languages. CodeQL uses queries to generate a relational database of multiple intermediate representations of the codebase, including the abstract syntax tree (AST), control flow graph (CFG), and data flow graph. For our baseline comparison, we utilized CodeQL queries (available in Github [2]) developed by Froh et al. We used their queries because 1) the queries were constructed researching npm malicious packages; 2) CodeQL is a mature commercial framework; and 3) a static analyzer as a pre-screener before LLM analysis is cost-efficient in a production environment.
We used 39 out of 40 queries developed by Froh et al., excluding the query, to detect the introduction of new dependencies. Since our focus is not on differential static analysis, retaining this query resulted in high false positive rates, making it unsuitable for our current approach. The 39 remaining queries include code injection (10) involving the use of functions like eval for executing dynamically generated code; sensitive data exposure (4) focusing on unauthorized access to sensitive user and system files; network issues (4) identifying unauthorized data flows potentially indicating exfiltration or backdoor attempts; file system access (3) monitoring access to files that may indicate security breaches; obfuscation and encoding (5) detecting code obfuscation techniques used to hide malicious intent; and miscellaneous (12) covering other concerns such as unsafe handling of domain names and cryptographic functions.
@@ IV-A2 ChatGPT Models
We leverage the SecurityAI workflow discussed in Section III to review 18,754 unique files using GPT-3 and GPT-4. We ran both models on each file and then collected the final reports discussed in Section III-D3. We used memoization [13] to store and reuse file processing results, ensuring no file was processed more than once. We also implemented extended TTL caching to maintain the validity of cached results over time, providing economies of scale. According to our evaluation criteria (Section III-D4), a malware score of 0.5 marks the threshold between ’possible malware behavior’ (0.25-0.5) and ’likely malicious behavior’ (0.5-0.75). We aimed to assess the accuracy of LLMs in generating these scores, considering a package malicious if any file within it scored above 0.5.
@@ IV-B2 GPT Model’s Performance
V RQ2: Effectiveness of SecurityAI
In this section, we discuss RQ2: What is the effectiveness of using a static analyzer as a pre-screener with SecurityAI in terms of the number of files that need to be analyzed and the associated cost?
@@ V-A RQ2: Methodology
An LLM workflow can be expensive if practitioners want to review at scale (in millions of \$ for GPT-4). Additionally, OpenAI has an API call rate limits [38]. Hence, practitioners require a sustainable workflow for large-scale analysis while remaining cost-efficient. Motivated by prior studies [37] in this study, we investigate the effectiveness of using static analyzers as a pre-screening step before our LLM workflow. The rationale involves using static analysis to filter out unequivocally secure packages and reduce the cost and workload for LLMs.
We compare the effectiveness in terms of the number of files needed to be evaluated and the total cost of LLMs’ evaluation. We designed the experiment in which we first scanned all files (a total of 18,754 files) using both GPT models and calculated the associated costs. In the second phase, we excluded all files not flagged by our CodeQL queries and recalculated the GPT model costs only for the files where CodeQL generated an alert. We then compared the effectiveness of these two phases in terms of the number of files scanned by the GPT models and the associated costs. Note that we did not conduct a detailed time analysis, as the response times for both GPT models were similar (approximately 24 days per model) and often influenced by OpenAI’s performance, including the need for re-runs due to technical issues. CodeQL took one day to run all the queries. Since we did not thoroughly evaluate false positives and negative cases of the static analyzer used in the prior study [16], comparing CodeQL with GPT run time did not provide a balanced assessment. Figure 2 shows our comparison of file and cost analysis.
@@ VI-B1 True Positive (TP)
In our analysis, both models detected malicious code in 2,057 packages out of 2,180 malicious packages. GPT-3 and GPT-4 independently detected 2,128 and 2,089 packages, respectively. Our manual evaluation by two researchers further revealed a significant concentration of malicious packages, primarily targeting data theft or exfiltration (721 cases) to harvest sensitive user information. The analysis also indicated 201 instances of reverse shells, facilitating unauthorized remote access; and 715 instances of hidden backdoors embedded within packages for system entry. Additionally, the models detected 571 cases of malicious connecting to suspicious domains, including Pastebin and domains controlled by attackers. Typosquatting was identified in 12 instances, exploiting user errors by registering misleading domain names to distribute malware. Dependency confusion attacks were detected in 151 cases, tricking package managers into downloading malicious versions of packages by exploiting naming similarities with legitimate dependencies.
@@
Data Theft or Exfiltration 721
Hidden Backdoors 715
Connecting to Suspicious Domains 571
Reverse Shells 201
Dependency Confusion 151
Typosquatting 12
@@
We observed models accurately detected obfuscated code in 1,260 packages. Additionally, we observed GPT-3 assigned malware scores above 0.5 to a significantly higher number of files (2,226 files) compared to GPT-4 (1,025 files), resulting in 1,191 additional files requiring manual review. Among these additional files, 94% of the files were package.json files- flagged due to suspicious package names, abnormal version numbers, installation scripts, descriptions, dependencies, and maintainer details. However, GPT-4 assigned more moderate malware scores (between 0.25 and 0.5) to such files, indicating a more accurate assessment. This mismatch in scoring and reasoning highlights the GPT-3’s inference based on incomplete evidence, leading to an unfaithful score relative to the actual malicious pattern. We tagged these packages as true positive because GPT-3 correctly identified actual malicious files as well. However, high malware scores for non-malicious files indicate GPT-3 frequently hallucinates, generating unfaithful content based on the provided data [24]. Further discussion on hallucinations can be found in Sections VI-B2 and VI-B3.
@@ VI-B2 False Positives (FP)
GPT-3 identified 195 neutral packages as malicious packages exhibiting signs of potential error in the pattern recognition capabilities of GPT-3. GPT-4’s had only 3 false positives, indicating a superior analytical capacity. Our manual analysis showed that GPT-3 flagged legitimate URL sanitizer code, use of post-installation scripts, dynamic code execution, and others as potential threats without adequately considering security vulnerabilities, code practices context, or legitimate use cases. We found 22 instances where GPT-3 incorrectly classified security vulnerabilities as malware, but GPT-4 accurately assigned a high-security risk but low malware score. Additionally, 173 packages were tagged as malicious by GPT-3 because of hallucination. Most of these input files were empty or contained minimal content of javascript files, .md, or .json files. For example, a readme.md file contain a text “place holder”, GPT-3 concluded that— The code is likely gathering system information and sending it to an external server. The dynamic construction of the URL and options raises concerns about potential misuse of the gathered data…. On the contrary, GPT 4 concluded — The provided content does not contain any JavaScript code to analyze; it’s just placeholder text. We suspect GPT-3 hallucinates because of a mismatch between expectations and insufficient information in the input. Such hallucination also highlights GPT-3’s limitations over GPT-4 in iterative self-refinement prompting (Section  III-A). Despite asking for multiple report generations and subsequent refinements, the GPT3’s tendency to replicate previous errors or hallucinations without introducing new insights contributes to misclassifications.
In the case of GPT-4, the model flagged a package.json file as suspicious due to the inclusion of a new dependency version, ’eslint-plugin-import’: ’npm:eslint-plugin-i@2.28.0-2’, mistaking it for a potential dependency confusion attack. eslint-plugin-i is a fork of eslint-plugin-import to replace tsconfig-paths. Though GPT-4 showed attention in this instance, which is beneficial for security, it may still struggle with new information after its last training cutoff in April 2023. Then, GPT-4 incorrectly raised concerns about a module http: 0.0.1-security due to an unusual version number, which was likely a developer’s error (trying to add a security holding npm packages as a dependency) rather than an intentional security threat.
@@ VI-B3 False Negative (FN)
Our study had 142 false negative cases where models did not identify the true malicious code. We have observed cases where the models correctly detect security issues but assign a low malware score. An exemplary case of GPT-4— The code exhibits potentially malicious behavior by collecting sensitive user information and sending it to a remote location under certain conditions… Despite these red flags, the model assigned a low-risk score of 0.4, suggesting a disconnect between the detection of individual risk factors and the aggregate assessment of malicious intent.
@@
VII Discussion: Challenges in using ChatGPT
VII-A Mode Collapse & Hallucination
@@
VII-B Large File & Intraprocedural Analysis
LLMs face challenges in analyzing large files due to several inherent limitations in token usage and contextual understanding. Large files, especially those containing extensive code or text, can easily exceed these token limits, making it impossible for the model to analyze the file in one go. Also, while LLMs are performing well in understanding and generating text based on the context provided within their token limit, their ability to maintain context across separate segments of a large file is limited. Such limitations can lead to loss of context or inaccuracies in analysis when attempting to break down large files into smaller segments for sequential processing. Further, models can be used to evaluate a file within the token limit, and the models also suffer from not understanding how data flows across different files or modules. Therefore, if malware or a vulnerability is spread across multiple files, the model could miss it because of limitations in processing the entire package scope.
VII-C Prompt Injection
@@
VII-D Parsing Output
We want SecurityAI’s output to be machine-readable. To do that, we ask models to generate JSON-format output and include an example of the expected report (Table I-D11). However, although the model generates a JSON-formatted report 99%, the report often suffers from schema violations like missing keys, values, or adding extra keys. Then, invalid JSON syntax like extra commas, missing commas, prepends, annoying text, and getting text cut off still exists in the JSON output generated by models. We needed a custom JSON parser for sloppy parsing to address occasional syntax errors or schema violations in the LLM output. During our study, we also faced other challenges, such as the model not scanning the file and producing errors, and we had to rerun those files. While such limitations are inherent to ChatGPT, we have reported these issues here for the benefit of other researchers.

### v3

@@ III-C3 Step 3: Final Report
The final report prompt (Table III) is designed to select and refine the best report from step 2. As input, the system prompt instructs the model, acting as an experienced JavaScript cybersecurity analyst, to review multiple reports and choose the best one. The task involves eliminating irrelevant signals by reviewing reports from step 2 and setting the malware score to 1 if the code is malicious to emphasize consistency between the final score and the report content. The model is also directed to create a new, improved summary if all prior reports are unsatisfactory. The step also involves sequentially feeding the model system role prompts from Step 1 (Table I) to remind the objectives and ensure a thorough code reevaluation. Overall, the model uses system role prompts (Table III and Table I), the input content (npm packages files), and reports from step 2 to output a conclusive summary report, operating at a lower temperature of 0.5 and a $top_{p}$ value of 0.5 to enhance focus and precision.
III-C4 Step 4: Report Evaluation
@@
In this step, a human evaluates the final report generated in step 3. Table IV contains the shell scripts of real-world malicious packages “corporate-delegate-packages” and the associated GPT-4-generated final report. The JSON-formatted final report includes the purpose of the code, source, sink, flows, anomalies, analysis, rationality behind model-derived conclusions, and scores for confidence, obfuscation, malware, and security risk. Reviewers can prioritize evaluations based on the malware score. Based on our scoring rules, a package is prioritized for human evaluation if any file has a malware score exceeding 0.5 (Section IV-A2). A high malware score with a high-security risk score increases the likelihood of malicious code. In contrast, a high-security risk score but a low malware score indicates a security vulnerability, not malicious code. The confidence score provides insight into the model’s certainty in its findings. An obfuscated score is helpful to understand whether the attacker used obfuscation to hide the code intent. Reviewers can review the purpose and conclusion parts (R1 & R7 in Table  IV) to comprehend the code’s intent. If further verification is needed, they can analyze the code review metrics (R2-R6 in Table IV).
@@ IV-A1 CodeQL Static Analysis
For our baseline comparison, we selected the static analyzer over any other existing technique to reduce the production cost of pre-screening. We chose not to adopt other existing static analyzers focused on malicious detection due to several factors: 1) a high false positive rate [35, 8], reliance on combined detection techniques such as static, dynamic, and machine learning [9], whereas we used static rules only to minimize overhead, and 3) the tools were tailored to detect malicious commits in VCS [14] or targeted languages like Python [29] and Java [24], rather than our focus on npm.
Froh et al., in a recent 2023 study [11], introduced a differential static analyzer leveraging custom CodeQL queries to identify JavaScript malicious code. For our baseline comparison, we utilized CodeQL queries (available in Github [2]) developed by Froh et al.  [11]. We utilized CodeQL over other existing static tools because 1) CodeQL is a commercially mature, state-of-the-art static analyzer, open source, and more developed than other static analyzers proposed in research for malicious package detection [33]; 2) the CodeQL queries used by Froh et al. [7] were explicitly designed for npm malicious packages, aligning with our dataset selection.
We used 39 out of 40 queries developed by Froh et al., excluding the query, to detect the introduction of new dependencies. Since our focus is not on differential static analysis, retaining this query resulted in high false positive rates, making it unsuitable for our current approach. The 39 remaining queries include code injection- involving the use of functions like eval for executing dynamically generated code; sensitive data exposure- focusing on unauthorized access to sensitive user and system files; network issues- identifying unauthorized data flows potentially indicating exfiltration or backdoor attempts; file system access- monitoring access to files that may indicate security breaches; obfuscation and encoding- detecting code obfuscation techniques used to hide malicious intent; and miscellaneous- covering other concerns such as unsafe handling of domain names.
@@ IV-A2 ChatGPT Models
We leverage the SocketAI workflow discussed in Section III to review 18,754 unique files using GPT-3 and GPT-4. We ran both models on each file and then collected the final reports discussed in Section III-C3. According to our evaluation criteria (Section III-C4), a malware score of 0.5 marks the threshold between ’possible malware behavior’ (0.25-0.5) and ’likely malicious behavior’ (0.5-0.75). We aimed to assess the accuracy of LLMs in generating these scores, considering a package malicious if any file within it scored above 0.5.
@@ IV-B2 GPT Model’s Performance
V RQ2: Effectiveness of SocketAI
In this section, we discuss RQ2: What is the effectiveness of using a static analyzer as a pre-screener with SocketAI in terms of the number of files that need to be analyzed and the associated cost?
@@ V-A RQ2: Methodology
An LLM workflow can be expensive if practitioners want to review at scale (in millions of \$ for GPT-4). Additionally, OpenAI has an API call rate limits [34]. Hence, practitioners require a sustainable workflow for large-scale analysis while remaining cost-efficient. Our study investigates the effectiveness of static analyzers as a pre-screening step before our LLM workflow. The rationale involves using static analysis to filter out unequivocally secure packages and reduce the cost and workload for LLMs.
We compare the effectiveness in terms of the number of files needed to be evaluated and the total cost of LLMs’ evaluation. We designed the experiment in which we first scanned all files (a total of 18,754 files) using both GPT models and calculated the associated costs. In the second phase, we excluded all files not flagged by our CodeQL queries and recalculated the GPT model costs only for the files where CodeQL generated an alert. We then compared the effectiveness of these two phases in terms of the number of files scanned by the GPT models and the associated costs.
Note that we did not conduct a detailed time analysis, as the response times for both GPT models were similar ( 24 days per GPT model and approximately 190 hours of qualitative study per independent researcher) and often influenced by OpenAI’s performance, including the need for re-runs due to technical issues. CodeQL took one day to run all the queries. Since we did not thoroughly evaluate false positives and negative cases of the static analyzer used in the prior study [11], comparing CodeQL with GPT run time did not provide a balanced assessment. Figure 2 shows our comparison of file and cost analysis.
@@ VI-B1 True Positive (TP)
In our analysis, the overlapping set of malicious packages detected by both models comprised 2,057 out of a total of 2,180 malicious packages. GPT-3 and GPT-4 independently detected 2,128 and 2,089 packages, respectively. Our manual evaluation by two researchers further revealed a significant concentration of malicious packages, primarily targeting data theft or exfiltration (736 cases) to harvest sensitive user information. The analysis also indicated 70 instances of reverse shells, facilitating unauthorized remote access, and 126 instances of hidden backdoors embedded within packages for system entry. Additionally, the models detected 391 cases of malicious connecting to suspicious domains, including Pastebin and domains controlled by attackers. Typosquatting was identified in 19 instances, exploiting user errors by registering misleading domain names to distribute malware. Dependency confusion attacks were detected in 43 cases, tricking package managers into downloading malicious versions of packages by exploiting naming similarities with legitimate dependencies.
@@
Data Theft or Exfiltration 736
Hidden Backdoors 126
Connecting to Suspicious Domains 391
Reverse Shells 70
Execution of Arbitrary Code 243
Dependency Confusion 43
Typosquatting 19
@@
We observed GPT-3 assigned malware scores above 0.5 to a significantly higher number of files (2,226 files) compared to GPT-4 (1,025 files), resulting in 1,191 additional files requiring manual review. Among these additional files, 94% of the files were package.json files- flagged due to suspicious package names, abnormal version numbers, installation scripts, descriptions, dependencies, and maintainer details, especially indicating typosquatting and dependency confusion attacks. However, typosquatting and dependency confusion requires correlating findings from both the package.json file and the source code to determine the attack’s type. In such cases, GPT-4 assigned more moderate malware scores (between 0.25 and 0.5) to such files and detected the actual issues in the source code, indicating a more accurate assessment. This mismatch in scoring and reasoning highlights the GPT-3’s inference based on incomplete evidence. We tagged these packages as true positive because GPT-3 also correctly identified actual malicious files. However, high malware scores for non-malicious files indicate GPT-3 frequently hallucinates, and additional files require manual review. Further discussion on hallucinations can be found in Sections VI-B2 and VI-B3.
@@ VI-B2 False Positives (FP)
GPT-3 identified 195 neutral packages as malicious packages exhibiting signs of potential error in the pattern recognition capabilities of GPT-3. Our manual analysis showed that GPT-3 flagged legitimate URL sanitizer code, use of post-installation scripts, dynamic code execution, and others as potential threats without adequately considering security vulnerabilities, code practices context, or legitimate use cases. We found 22 instances where GPT-3 incorrectly classified security vulnerabilities as malware, but GPT-4 accurately assigned a high-security risk but low malware score. Additionally, 173 packages were tagged as malicious by GPT-3 because of hallucination. Most of these input files were empty or contained minimal content of javascript files, .md, or .json files. For example, a readme.md file contain a text “place holder”, GPT-3 concluded that— The code is likely gathering system information and sending it to an external server. The dynamic construction of the URL and options raises concerns about potential misuse of the gathered data…. On the contrary, GPT 4 concluded — The provided content does not contain any JavaScript code to analyze; it’s just placeholder text. We suspect GPT-3 hallucinates because of a mismatch between expectations and insufficient information in the input. Such hallucination also highlights GPT-3’s limitations over GPT-4 in iterative self-refinement prompting (Section  III-A). Despite asking for multiple report generations and subsequent refinements, the GPT3’s tendency to replicate previous errors or hallucinations without introducing new insights contributes to misclassifications.
GPT-4’s had only 3 false positives. In one of the cases, the model flagged a package.json file as suspicious due to the inclusion of a new dependency version, ’eslint-plugin-import’: ’npm:eslint-plugin-i@2.28.0-2’, mistaking it for a potential dependency confusion attack. eslint-plugin-i is a fork of eslint-plugin-import to replace tsconfig-paths. Though GPT-4 showed attention in this instance, which is beneficial for security, it may still struggle with new information after its last training cutoff in April 2023. Then, GPT-4 incorrectly raised concerns about a module http: 0.0.1-security due to an unusual version number, which was likely a developer’s error (trying to add a security holding npm packages as a dependency) rather than an intentional security threat.
@@ VI-B3 False Negative (FN)
Our study had 142 false negative cases where both models did not identify the true malicious code. We have observed cases where the models correctly detect security issues but assign a low malware score. An exemplary case of GPT-4— The code exhibits potentially malicious behavior by collecting sensitive user information and sending it to a remote location under certain conditions… Despite these red flags, the model assigned a low-risk score of 0.4, suggesting a disconnect between the detection of individual risk factors and the aggregate assessment of malicious intent.
@@
VII Discussion
In this section, we delve into the broader implications of using SocketAI and explore the challenges and threats to validity encountered in this study.
VII-A Implications of SocketAI and future work:
Our contribution paves the way for new tool development and future research in AI-driven software development and software security domain. While prior studies use LLMs to detect vulnerabilities [36], to our knowledge, ours is the first to propose an end-to-end LLM-based workflow for malicious code review. Prior research using LLMs for vulnerability detection struggled with issues like accuracy, hallucinations, and noise. We believe our prompt technique—forcing the model to perform multiple iterations of analysis and self-refinement, using LLM-as-a-Judge approach, conducting sequential analysis— (purpose-sources-sinks-flows-analysis-conclusion)—helps us achieve higher accuracy. Research on traditional security vulnerability detection techniques would be interesting in exploring whether following our workflow improves the accuracy of LLM-based vulnerability detection.
Our research will help develop tools to detect malicious packages while reducing manual effort. Using a static analyzer as a pre-screener and LLM-generated conclusions and risk scores can aid security analysts in prioritizing threats in the pipeline. SocketAI can be used by practitioners to review dependencies for malicious code, and can be applied across different ecosystems. Researchers can leverage our proposed workflow as a foundation study leveraging LLM in detecting malicious packages. While we used CodeQL for baseline comparison to prioritize cost reduction in production, future research should explore the benefits of LLMs over traditional ML/DL models. Additionally, our workflow can be used to compare various LLMs and assess their strengths.
VII-B Challenges in using ChatGPT
VII-B1 Mode Collapse & Hallucination
@@
VII-B2 Large File & Intraprocedural Analysis
LLMs face challenges in analyzing large files due to several inherent limitations in token usage and contextual understanding. Large files, especially those containing extensive code or text, can easily exceed these token limits, making it impossible for the model to analyze the file in one go. Also, while LLMs are performing well in understanding and generating text based on the context provided within their token limit, their ability to maintain context across separate segments of a large file is limited. Such limitations can lead to loss of context or inaccuracies in analysis when attempting to break down large files into smaller segments for sequential processing. Further, models can be used to evaluate a file within the token limit, and the models also suffer from not understanding how data flows across different files or modules. Therefore, if malware (e.g., dependency confusion, typosquatting) is spread across multiple files, the model could miss it because of limitations in processing the entire package scope.
VII-B3 Prompt Injection
@@
VII-B4 Parsing Output
We want SocketAI’s output to be machine-readable. To do that, we ask models to generate JSON-format output and include an example of the expected report (Table I-D11). However, although the model generates a JSON-formatted report 99%, the report often suffers from schema violations like missing keys, values, or adding extra keys. Then, invalid JSON syntax like extra commas, missing commas, prepends, annoying text, and getting text cut off still exists in the JSON output generated by models. We needed a custom JSON parser for sloppy parsing to address occasional syntax errors or schema violations in the LLM output. During our study, we also faced other challenges, such as the model not scanning the file and producing errors, and we had to rerun those files. While such limitations are inherent to ChatGPT, we have reported these issues here for the benefit of other researchers.
VII-C Threats To Validity
MalwareBench dataset [55] may contain bias toward a certain type of malicious code, and the repetition of the same attack across multiple packages might have skewed the dataset. As discussed in Section VII-B, model performance might differ with complex attacks or large files where malware spreads in different forms all over the file. Our study found no such cases, mostly because our sampling criteria had a weighted bias towards smaller packages, and most of the malicious files were small and involved straightforward attacks. Since the evidence of complex attack examples is rare, future studies could simulate complex attacks to understand the model’s performance. Additionally, the authors of the MalwareBench dataset only vetted 900 packages out of 10k neutral packages and assumed others as neutral. Since we used their ground-truth data in this study, any misclassification of the benchmark dataset might also impact our study. We did our best to check that our dataset’s neutral samples do not include malicious code by manually evaluating 2,530 files out of 20,926 files. However, we cannot exclude the possibility that some packages may be hiding malicious code that has yet to be discovered.
Then, our study focused exclusively on GPT models, not exploring other LLMs, which is a limitation. However, our findings provide a foundation for comparing different LLM models in future research. Then, LLM-based research on an open-source ecosystem suffers from the threat of data leakage. To mitigate this, we used zero-shot learning and selected packages from MalwareBench (close source and access given upon request) that contain packages released before and after the training cut-off dates (April 2023 for GPT-4 and September 2021 for GPT-3). Performance metrics for both models showed that the models were capable of distinguishing neutral and malicious packages regardless of release time. Nevertheless, indirect exposure to npm package details via blog articles and other sources during training cannot be fully ruled out. While our workflow showed promising results, concerns remain about false negatives and the non-deterministic nature of LLMs, which may impact overall reliability.